### v2

@@ Abstract
Blue straggler stars in stellar clusters are a subset of stars that are bluer and appear younger than other cluster members, seemingly straggling behind in their evolution. They offer a unique opportunity to understand the stellar dynamics and populations within their hosts. In the collisional formation scenario, a persistent challenge is the excessive angular momentum in the collision product. The consequent significant mass loss during transition to a stable state leads to a star with too low a mass to be a blue straggler, unless it spins down efficiently. While many proposed spin-down mechanisms involve boosted angular momentum loss via magnetic braking within the collision product, the existence or strength of these magnetic fields has not been confirmed. Here, we report three-dimensional magnetohydrodynamical simulations of collisions between two low-mass main-sequence stars and investigate magnetic field amplification. Magnetic field energy is amplified by a factor of $10^{8}-10^{10}$, resulting in the magnetic field strength of $10^{7}-10^{8}$ at the core of the collision product, independent of collision parameters. The surface magnetic field strengths have increased up to $10-10^{4}$ G. In addition, a distinctly flattened, rotating gas structure appears around the collision products in off-axis collisions, which may be a hint of possible disk formation. Such significant magnetic amplification and potential disk formation suggest the possibility of efficient spin-down of collision products via magnetic braking and magnetic disk locking, which can result in their appearance as blue stragglers.
keywords:
Magnetohydrodynamics, Stellar collision, globular cluster, blue straggler
1 Main
@@
Stars can collide physically in dense stellar environments, such as globular cluster cores. For typical stellar densities of globular clusters, up to tens of percent of stars in the center can undergo a single collision in their lifetime [1, 2]. These collisions are not sufficiently energetic, so the two stars merge into a more massive star, rather than both stars being destroyed. Because of chemical mixing, heat injection, and rotation induced by collision, the collision products can manifest as blue stragglers [3, 4]–a group of stars that appear bluer and younger than their counterparts and are located differently from most cluster members on color-magnitude diagrams [1, 5]. Stellar collisions, facilitated by multi-body interactions involving binaries [6, 7, 8], may serve as the dominant formation channel for blue stragglers in dense clusters [9]. These stars can provide unique information about the evolution history of clusters and stellar populations within them.
@@
The properties of collision products have been studied using hydrodynamics simulations [e.g., 10, 11, 12] and their subsequent long-term evolution using stellar evolution simulations [e.g., 13, 5, 14]. These studies have shown that collision products are out of thermal equilibrium upon collision and, therefore, they are swollen in size. In addition, unless the collision is perfectly head-on, the collision product contains angular momentum comparable to or possibly larger than the theoretical upper limit of the angular momentum for a normal star of the same mass [11, 12, 15]. This naturally leads the collision product to lose a significant fraction of its mass and angular momentum during transition to a thermal equilibrium state [11, 12], unless it spins down efficiently. This excessive angular momentum of the collision product posed an ‘angular momentum’ problem for the formation of blue stragglers [12], where if the collision product loses too much mass before it spins down, the outcome would have too low mass to appear as blue stragglers.
@@
Consequently, whether collisions are a plausible mechanism for the formation of blue stragglers comes down to a key question: can the collision product spin down efficiently? While a few spin-down mechanisms have been proposed, most of them require the existence of magnetic fields within the star, such as magnetic braking [16, 17, 18]. In addition to the direct impact of magnetic fields on the spin, it has been known that magnetic fields can also significantly affect the internal structure of the star, namely, the redistribution of angular momentum and chemical elements inside rotating stars [19], and, therefore, even their final fate. Despite the importance of magnetic fields, the magnetic field strength and configuration inside collision products remain highly uncertain and no work on this subject using magnetohydrodynamics simulations in the context of blue stragglers has been done in the last decade.
In this article, we investigate the amplification of magnetic fields during collisions between realistic low-mass main-sequence stars, their field configuration and internal structure, using moving-mesh magnetohydrodynamics simulations.
@@ 2 Results
We performed simulations of parabolic [^1] This assumption for the orbit is reasonable, considering that the typical eccentricity for collisions in globular clusters with velocity dispersion $\sigma\simeq 10-15$ km s-1 [20] can be estimated as $|1-e|\simeq 3\times 10^{-4}(\sigma/10{\rm km~{}s}^{-1})^{2}(M/2\,\mathrm{M}_{\odot})^{-1}(r_{\rm p}/1\,\mathrm{R}_{\odot})$, where $M$ is the total mass of two colliding objects. For this estimate, we assume that the relative specific kinetic energy $\simeq 0.5\sigma^{2}$ and the specific angular momentum is $\simeq\sqrt{2GMr_{\rm p}}$. collisions between main-sequence stars such as those found in globular clusters, using the moving-mesh magnetohydrodynamics code AREPO [21, 22, 23]. We consider mass ratios between the two colliding stars ranging from 0.5 to 0.97, with primary masses close the main-sequence turn-off mass for globular clusters at 12 Gyr [24] ($M_{1}=0.7$ and $0.8\,\mathrm{M}_{\odot}$). The choice of the primary masses can be justified, as more massive stars tend to accumulate near the core, where collisions are most frequent, due to mass segregation. The impact parameter $b=r_{\rm p}/(R_{1}+R_{2})$ ranges from $0.1$ to $0.5$, covering from nearly head-on collisions ($b=0.1$) to strong grazing encounters ($b=0.5$). Here, $r_{\rm p}$ is the pericenter distance and $R_{1}$($R_{2}$) is the radius of the primary (secondary) star. See Table 1 for the complete list of our models with the initial collision parameters.
We assume the magnetic field of the original stars is described by a dipole field with a surface magnetic field strength of $B_{\rm s}=1$G, which is comparable to or smaller than the average magnetic field strength of the Sun [25, 26, 27], and smaller than that of M-dwarf stars [28]. The resolution is such that there are $100-125$ cells per stellar radius of the original star, corresponding to the mass resolution of $5\times 10^{-7}$ of the stellar mass on average and the numerical Reynolds number (Equation 2 in [29]) is $\simeq 1200-3000$, assuming that the turbulent injection scale $\simeq 0.5(R_{1}+R_{2})$. See §A for more details of our methods.
@@ 2.1 Magnetic field amplification and saturation
The magnetic field energy $E_{\rm B}$ has increased by a factor of $10^{8}-10^{10}$, reaching $E_{\rm B}\simeq 10^{44}-10^{45}$ erg, by the time the collision product has recovered a dynamically steady state, independent of the masses of the colliding stars, impact parameters, and initial field strength and orientation. As a representative case, Fig. 1 illustrates the evolution of the magnetic field energy, along with the density, passive scalar, and magnetic field strength distributions, for the collision between $0.7\,\mathrm{M}_{\odot}$ and $0.6\,\mathrm{M}_{\odot}$ stars with $b=0.5$. The passive scalar is an artificial scalar quantity initially assigned to each cell, which then evolves through advection but does not influence the evolution of hydrodynamics quantities. In this case, the two stars touch three times (at $t\lesssim 13$ hours), after which the less massive star is destroyed and mixed into the more massive star ($t\simeq 13-20$ hours). The merger product enters a relaxation phase ($t\gtrsim 20$ hours), the magnetic energy has reached $\simeq 10^{45}$ erg and the growth rate has dropped substantially. See the left panel of Fig. 2 for collisions between the same stars with different impact parameters and Fig. 6 for the rest of the models.
Figure 2: (Left) Evolution of the total magnetic field energy $E_{\rm B}$ for collisions between $0.7\,\mathrm{M}_{\odot}$ and $0.6\,\mathrm{M}_{\odot}$ stars, as representative cases, with three different impact parameters $b$, as a function of time since the first contact. The growth rate varies, depending on the number of contacts before the final merger. However, in all cases, $E_{\rm B}$ grows by more than eight orders of magnitude over a few tens of hours via small-scale dynamos, and then saturates at $10^{44}-10^{45}$erg or rises substantially more slowly. (Right) Radial magnetic field strength of the products, mass-weighted averaged over the spherical equipotential surfaces, as a function of the enclosed mass. The thinner lines indicate the dipole component defined as $\sqrt{(B^{\rm r})^{2}+(B^{\rm z})^{2}}$.
As illustrated by the representative case in Fig. 1, the collision process can be generally split into three stages, 1) contact, 2) merger, and 3) dynamical relaxation. In these collisions, there can be multiple moments of contact between the two stars before they finally merge. At each of these close approaches, turbulent and shearing flows amplify magnetic fields. This naturally leads to multiple episodes of sudden increases in $E_{\rm B}$ until it saturates, as illustrated in Figs. 1 and 2. This is clearly different from the magnetic field amplification of inspiral binary mergers [30, 31] where $E_{\rm B}$ increases relatively smoothly. The number of close encounters depends primarily on the impact parameter. For nearly head-on collisions (e.g., $b=0.1$), the two stars merge almost immediately after the first contact, during which $E_{\rm B}$ increases exponentially. However, for off-axis collisions (e.g., $b\gtrsim 0.25$), they encounter more than once before the final merger, during which the growth of $E_{\rm B}$ is not well-described by a single exponential or power-law curve.
@@
By the time the collision product enters the relaxation phase, the total magnetic field energy saturates at a similar level ($10^{44}-10^{45}$erg) in all our models, with a tendency of being distributed more towards the mid-plane. The small-scale dynamo theory predicts that the magnetic energy saturates when it becomes comparable to turbulent kinetic energy [32, 33, 28, 34, 35], which is consistent with our simulations (see Fig. 7). We also confirmed that the saturation level is robust against the resolution and initial $B$ strength, which further supports that the small-scale dynamo would be responsible for the magnetic field amplification.
@@
The magnetic field energy has saturated or grows noticeably more slowly in the dynamically relaxed products. By that time, most of the magnetic fields also are expected to have been relaxed. Almost all $E_{\rm B}$ is stored within 50% (90%) of the total mass of the collision products for $b\leq 0.25$ (0.5), where the local Alfvénic timescales are found to be shorter than the evolution time (see Fig. 8). Despite the dramatic amplification, the magnetic field energy comprises only a small fraction ($\lesssim 10^{-4}$) of the total energy (see Fig. 7). As expected, the magnetic pressure does not exceed the gas + radiation pressure across the product -- the ratio of the gas+radiation pressure to the magnetic pressure, also known as the plasma $\beta$, $\simeq 10$ at the core and $>10$ at the envelope.
@@
The dominant component of $E_{\rm B}$ (as well as the kinetic energy) within the collision products is the azimuthal component (see Fig. 7), with the saturated magnetic field strength $B$ near the center $\simeq 10^{7}-10^{8}$ G, as shown in the right panel of Fig. 2. In general, the gradient of $B$, relative to the enclosed mass, is shallower for higher impact parameters ($b$) -- the surface $B$ is a few G for $b=0.1$, while it is $\simeq 10^{3}$ G for $b=0.5$. The mass-weighted average of magnetic flux is $\simeq 10^{26}-10^{28}$ G cm2 and surface magnetic flux of $10^{23}-10^{26}$ G cm2. If the magnetic flux is conserved, the surface magnetic field of the thermally relaxed collision product would be $10-10^{4}$. The dipole field ($\sqrt{(B^{r})^{2}+(B^{z})^{2}}$, where $B^{r}$ is the radial component and $B^{z}$ is the vertical component), more relevant for the magnetic braking and, thus, the spin down of the merger product, is consistently smaller than the total field by a factor of a few.
@@
Figure 3: Mass-weighted average of the temperature $T_{\rm mid}$ in the mid-plane (left) and the ratio of the rotational velocity in the mid-plane to the Keplerian velocity $v_{\rm mid}^{\phi}/v^{\phi}_{\rm Kep}$, as a function of distance from the center of mass of the collision products in collisions involving $M_{1}=0.7\,\mathrm{M}_{\odot}$, with different impact parameters $b$. The temperature profile shows a noticeable slope change at $0.7-0.8\,\mathrm{R}_{\odot}$ for $b\leq 0.25$ (except for one case with $M_{2}=0.4\,\mathrm{M}_{\odot}$ and $b=0.25$) and $1-2\,\mathrm{R}_{\odot}$ for $b=0.5$ (solid lines). The velocity ratio profiles show the similar features.

### v3

@@ Abstract
Blue straggler stars in stellar clusters appear younger and bluer than other cluster members, offering a unique opportunity to understand the stellar dynamics and populations within their hosts. In the collisional formation scenario, excessive angular momentum in the collision product poses a challenge, as the consequent significant mass loss during transition to a stable state leads to a star with too low a mass to be a blue straggler, unless it spins down efficiently. While many proposed spin-down mechanisms require magnetic fields, the existence or strength of these magnetic fields has not been confirmed. Here, we present three-dimensional moving-mesh magnetohydrodynamical simulations of collisions between low-mass main-sequence stars and investigate magnetic field amplification. Magnetic field energy is amplified during collisions by a factor of $10^{8}-10^{10}$, resulting in the magnetic field strength of $10^{7}-10^{8}$ at the core of the collision product, independent of collision parameters. The surface magnetic field strengths increase up to $10-10^{4}$ G. In addition, a distinctly flattened, rotating gas structure appears around the collision products in off-axis collisions, suggesting potential disk formation. These findings indicate that magnetic braking and disk locking could facilitate spin-down, enabling the formation of blue straggler stars.
Magnetohydrodynamics -- Stellar collision -- globular cluster -- blue straggler
journal: ApJLsoftware: AREPO (Springel, 2010; Weinberger et al., 2020; Pakmor et al., 2016), MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019), Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020)
1 Introduction
@@
Stars can collide physically in dense stellar environments, such as globular cluster cores. For typical stellar densities of globular clusters, up to tens of percent of stars in the center can undergo a single collision in their lifetime (Hills & Day, 1976; Dale & Davies, 2006). These collisions are not particularly energetic, so the two stars merge into a more massive star, rather than both stars being destroyed. Due to chemical mixing, heat injection, and rotation induced by collisions, collision products can manifest themselves as blue stragglers (Sandage, 1953; Burbidge & Sandage, 1958)–a group of stars that appear bluer and younger than their counterparts and are located differently from most cluster members on color-magnitude diagrams (Hills & Day, 1976; Sills et al., 2009). Stellar collisions, facilitated by multibody interactions involving binaries (Leonard, 1989; Leonard & Fahlman, 1991; Leonard & Linnell, 1992), may serve as the dominant formation channel for blue stragglers in dense clusters (Chatterjee et al., 2013). These stars can provide unique information about the evolution history of clusters and stellar populations within them. See Wang & Ryu (2024) the most recent review for blue straggler stars.
@@
The properties of collision products have been studied using hydrodynamics simulations (e.g., Hills, 1988; Sills et al., 2001, 2005) and their subsequent long-term evolution using stellar evolution simulations (e.g., Glebbeek et al., 2008; Sills et al., 2009; Glebbeek et al., 2013). These studies have shown that collision products are out of thermal equilibrium upon collision and, therefore, they are swollen in size. In addition, unless the collision is perfectly head-on, the collision product contains angular momentum comparable to or larger than the theoretical upper limit of the angular momentum for a normal star of the same mass (Sills et al., 2001, 2005; Ryu et al., 2024b). This naturally leads the collision product to lose a significant fraction of its mass and angular momentum during the transition to a thermal equilibrium state (Sills et al., 2001, 2005), unless it spins down efficiently. This excessive angular momentum of the collision product posed an ‘angular momentum’ problem for the formation of blue stragglers (Sills et al., 2005), where if the collision product loses too much mass before it spins down, the outcome would have too low mass to appear as blue stragglers.
@@
Consequently, whether collisions are a plausible mechanism for the formation of blue stragglers comes down to a key question: Can the collision product spin down efficiently? While a few spin-down mechanisms have been proposed, most of them require the existence of magnetic fields within the star, such as magnetic braking (Weber & Davis, 1967; Leonard & Livio, 1995; Eggenberger et al., 2005). In addition to the direct impact of magnetic fields on the spin, it has been known that magnetic fields can also significantly affect the internal structure of the star, namely, the redistribution of angular momentum and chemical elements inside rotating stars (Takahashi & Langer, 2021), and, therefore, even their final fate. Despite the importance of magnetic fields, the magnetic field strength and configuration inside collision products remain highly uncertain, and no work on this subject using magnetohydrodynamic simulations in the context of blue stragglers has been done in the last decade.
In this letter, we investigate the evolution of magnetic fields in stellar collisions using moving-mesh magnetohydrodynamic simulations. Considering a wide range of parameters, we demonstrate that magnetic field energy is amplified significantly during collisions, independent of collision parameters. Additionally, we find evidence of disk formation around the collision products, especially when the impact parameter is large.
The paper is organized as follows. In § 2, we describe our methods for creating stellar models (§ 2.1) and magnetohydrodynamics simulations (§ 2.2), and provide the parameters considered (§ 2.3). We present our results in § 3. Finally, we summarize and discuss the implications of our results in § 4.
2 Methods
2.1 Initial stellar model
We take the initial state of main-sequence stars, computed using the stellar evolution code MESA  (version r24.03.1 Paxton et al., 2011, 2013, 2015, 2019; Jermyn et al., 2023). We adopt the prescriptions for convection, overshoot, semiconvection, and thermohaline mixing from Choi et al. (2016). We use the Ledoux criterion (Ledoux, 1947) to identify the boundary of the convective regions. We consider a range of stellar masses from $0.4-0.8\,\mathrm{M}_{\odot}$ at the stellar age of $12$ Gyr and a metallicity $Z=0.001$, relevant for old stars in globular clusters (e.g. VandenBerg et al., 2013). We show the density and H mass fraction profiles of our stellar models in Fig. 6 in Appendix A.
2.2 3D Magnetohydrodynamic simulations
We perform a suite of 3D magnetohydrodynamic (MHD) simulations of stellar collisions using the massively parallel gravity and MHD moving-mesh code AREPO (Springel, 2010; Weinberger et al., 2020; Pakmor et al., 2016). By adopting the novel approach of a moving mesh, AREPO inherits advantages of Eulerian finite-volume methods and Lagrangian particle methods, resulting in accurate shock capturing without introducing an artificial viscosity, small advection errors, and an efficient tracking of supersonic flows. The code has been used to study a variety of astrophysics problems, ranging from cosmology (e.g., Vogelsberger et al., 2014; Springel et al., 2018), tidal disruption events (e.g., Vynatheya et al., 2023; Xin et al., 2023; Ryu et al., 2024b, c), to common envelope (e.g. Ohlmann et al., 2016), binary mergers (e.g., Pakmor et al., 2012; Schneider et al., 2019; Pakmor et al., 2024a), and stellar collisions (Ryu et al., 2024a). We adopt the OPAL equation of state (Rogers & Nayfonov, 2002) and follow the advection of five isotopes ($\mathrm{H}$, ${}^{4}\mathrm{He}$, ${}^{12}\mathrm{C}$, ${}^{14}\mathrm{N}$, ${}^{16}\mathrm{O}$). We employ the Powell scheme (Powell et al., 1999) for $B-$field divergence control. For implementation details in AREPO, and code tests, we refer to Pakmor et al. (2011); Pakmor & Springel (2013).
Our MHD simulations of stellar collisions involve two steps. In the first step, we create a 3D MS star for each MESA model using the mesh construction method by Ohlmann et al. (2017). We evolve the stars in isolation until they are fully relaxed, which takes less than 10 $t_{\rm dyn}$. Here, the stellar dynamical time is defined as $t_{\rm dyn}=\sqrt{R_{\star}^{3}/GM_{\star}}$ where $M_{\star}$ ($R_{\star}$) is the stellar mass (radius). For the stars considered, $t_{\rm dyn}=O(10^{2}-10^{3})$ s. During this relaxation phase, we ignore the magnetic field inside the stars because the seed magnetic field (1G at the stellar surface) is so small that the magnetic pressure is completely negligible compared to the thermal (gas + radiation) pressure.
where $B_{\rm s}$ is the stellar surface field strength at the pole, $\hat{\textbf{r}}$ the unit position vector inside the star, and $\hat{\textbf{b}}$ the unit dipole vector.
Table 1: Model list and initial parameters, and properties of dynamically relaxed collision products. The initial collision parameters are provided up to the fifth column: mass of the original primary star $M_{1}[\,\mathrm{M}_{\odot}]$, mass of the original secondary star $M_{2}[\,\mathrm{M}_{\odot}]$, mass ratio $M_{2}/M_{1}$, and pericenter distance $b=r_{\rm p}/[R_{1}+R_{2}]$.
- $M_{1}$ $M_{2}$ $q$ $b$
1 0.7 0.68 0.97 0.10
2 0.7 0.68 0.97 0.25
3 0.7 0.68 0.97 0.50
4 0.7 0.60 0.86 0.10
5 0.7 0.60 0.86 0.25
6 0.7 0.60 0.86 0.50
7 0.7 0.40 0.57 0.10
8 0.7 0.40 0.57 0.25
9 0.7 0.40 0.57 0.50
10 0.8 0.70 0.87 0.10
11 0.8 0.70 0.87 0.25
12 0.8 0.60 0.75 0.10
13 0.8 0.60 0.75 0.25
14 0.8 0.40 0.50 0.10
15 0.8 0.40 0.50 0.25
@@ 2.1 Initial stellar model
2.3 Parameters
We perform simulations of parabolic [^1] This assumption for the orbit is reasonable, considering that the typical eccentricity for collisions in globular clusters with velocity dispersion $\sigma\simeq 10-15$ km s-1 (Cohen, 1983) can be estimated as $|1-e|\simeq 3\times 10^{-4}(\sigma/10{\rm km~{}s}^{-1})^{2}(M/2\,\mathrm{M}_{\odot})^{-1}(r_{\rm p}/1\,\mathrm{R}_{\odot})$, where $M$ is the total mass of two colliding objects. For this estimate, we assume that the relative specific kinetic energy $\simeq 0.5\sigma^{2}$ and the specific angular momentum is $\simeq\sqrt{2GMr_{\rm p}}$. collisions between main-sequence stars such as those found in globular clusters. We consider mass ratios between the two colliding stars ranging from 0.5 to 0.97, with primary masses close the main-sequence turn-off mass for globular clusters at 12 Gyr (VandenBerg et al., 2013) ($M_{1}=0.7$ and $0.8\,\mathrm{M}_{\odot}$). The choice of the primary masses can be justified, as more massive stars tend to accumulate near the cluster core, where collisions are most frequent, due to mass segregation. The impact parameter $b=r_{\rm p}/(R_{1}+R_{2})$ ranges from $0.1$ to $0.5$, covering from nearly head-on collisions ($b=0.1$) to strong grazing encounters ($b=0.5$). Here, $r_{\rm p}$ is the pericenter distance and $R_{1}$($R_{2}$) is the radius of the primary (secondary) star. See Table 1 for the complete list of our models with the initial collision parameters.
We assume the magnetic field of the original stars is described by a dipole field with a surface magnetic field strength of $B_{\rm s}=1$G, which is comparable to or smaller than the average magnetic field strength of the Sun (Hale, 1913; Babcock & Babcock, 1955; Alfvén, 1956), and smaller than that of M-dwarf stars (Reiners et al., 2022). As Equation 1 indicates, the magnetic field increases inward inside the star. To be conservative, we limit the seed magnetic field to be less than $100B_{\rm s}$.
The resolution is such that there are $100-125$ cells per stellar radius of the original star, corresponding to the mass resolution of $5\times 10^{-7}$ of the stellar mass on average and the numerical Reynolds number (Equation 2 in Pakmor et al., 2024b) is $\simeq 1200-3000$, assuming that the turbulent injection scale $\simeq 0.5(R_{1}+R_{2})$.
We follow the evolution until a collision product forms and becomes fully relaxed dynamically whose internal structure (e.g., density, angular frequency, and magnetic field strength) does not change over time. The timescale of the collision process, from the initial contact to dynamical relaxation, is shorter for a smaller impact parameter. It takes around 20 - 50 hours since collision for $b\leq 0.25$, $>50$ hours for $b=0.5$. In all simulations, the fractional errors in total energy remain small ($\lesssim$ a few %) until the end of the simulations.
To ensure the robustness of our simulations, we performed resolution tests as well as additional simulations with different initial magnetic field strengths. We found that while the resolution and initial magnetic field strengths affect magnetic field amplification rates, the total saturated magnetic field energy inside collision products consistently reached a similar level. We refer to Appendix B for more details.
@@ 3 Results
3.1 Overview
The collision process can be generally split into three stages, 1) contact, 2) merger, and 3) dynamical relaxation. As a representative case, Fig. 1 illustrates the evolution of the magnetic field energy, along with the density, passive scalar, and magnetic field strength distributions, for the collision between $0.7\,\mathrm{M}_{\odot}$ and $0.6\,\mathrm{M}_{\odot}$ stars with $b=0.5$. The passive scalar is an artificial scalar quantity initially assigned to each cell, which then evolves through advection but does not influence the evolution of hydrodynamics quantities. In this case, the two stars touch three times (at $t\lesssim 13$ hours), after which the less massive star is destroyed and mixed into the more massive star ($t\simeq 13-20$ hours). The collision product enters a relaxation phase ($t\gtrsim 20$ hours), the magnetic energy has reached $\simeq 10^{45}$ erg and the growth rate has dropped substantially.
In collisions, there can be multiple moments of contact between the two stars before they finally merge. At each of these close approaches, turbulent and shearing flows amplify magnetic fields. This naturally leads to multiple episodes of sudden increases in $E_{\rm B}$ until it saturates. This is clearly different from the magnetic field amplification of inspiral binary mergers (Schneider et al., 2019; Kiuchi et al., 2024) where $E_{\rm B}$ increases relatively smoothly. The number of close encounters depends primarily on the impact parameter. For nearly head-on collisions (e.g., $b=0.1$), the two stars merge almost immediately after the first contact, during which $E_{\rm B}$ increases exponentially. However, for off-axis collisions (e.g., $b\gtrsim 0.25$), they encounter more than once before the final merger, during which the growth of $E_{\rm B}$ is not well-described by a single exponential or power-law curve.
@@ 3.2 Magnetic field amplification and saturation
The magnetic field energy $E_{\rm B}$ has increased by a factor of $10^{8}-10^{10}$, reaching $E_{\rm B}\simeq 10^{44}-10^{45}$ erg, by the time the collision product has recovered a dynamically steady state, independent of the masses of the colliding stars, impact parameters, and initial field strength and orientation, which is demonstrated in Fig. 2. By the time the collision product enters the relaxation phase, the total magnetic field energy saturates at a similar level in all our models, with a tendency of being distributed more towards the mid-plane. The small-scale dynamo theory predicts that the magnetic energy saturates when it becomes comparable to turbulent kinetic energy (Cattaneo, 1999; Federrath, 2016; Reiners et al., 2022; Kriel et al., 2023; Beattie et al., 2023), which is consistent with our simulations (see Fig. 8 in Appendix C). We also confirmed that the saturation level is robust against the resolution and initial $B$ strength, which further supports that the small-scale dynamo would be responsible for the magnetic field amplification.
Figure 2: Evolution of the total magnetic field energy $E_{\rm B}$ in all our models, as a function of time since the first contact. The growth rate varies, depending on the number of contacts before the final merger. However, in all cases, $E_{\rm B}$ grows by more than eight orders of magnitude over a few tens of hours via small-scale dynamos, and then saturates at $10^{44}-10^{45}$erg or rises substantially more slowly.
@@
The magnetic field energy has saturated or grows noticeably more slowly in the dynamically relaxed products. By that time, most of the magnetic fields also are expected to have been relaxed. Almost all $E_{\rm B}$ is stored within 50% (90%) of the total mass of the collision products for $b\leq 0.25$ (0.5), where the local Alfvénic timescales are found to be shorter than the evolution time.
Despite the dramatic amplification, the magnetic field energy comprises only a small fraction ($\lesssim 10^{-4}$) of the total energy (see Fig. 8). As expected, the magnetic pressure does not exceed the gas + radiation pressure across the product -- the ratio of the gas+radiation pressure to the magnetic pressure, also known as the plasma $\beta$, $\simeq 10$ at the core and $>10$ at the envelope. As a consequence, the overall hydrodynamic properties (e.g., density, pressure, and angular momentum profiles) and chemical element profiles inside the collision products are not significantly affected by the existence of magnetic fields.
Figure 3: Radial magnetic field strength of the products, mass-weighted averaged over the spherical equipotential surfaces of collision products produced in collisions between $0.7\,\mathrm{M}_{\odot}$ and $0.6\,\mathrm{M}_{\odot}$ stars with different impact parameters $b$, as a function of the enclosed mass. The thinner lines indicate the dipole component defined as $\sqrt{(B^{\rm r})^{2}+(B^{\rm z})^{2}}$, where $B^{\rm r}(B^{\rm z})$ is the radial (vertical) component of magnetic fields.
@@
The dominant component of $E_{\rm B}$ (as well as the kinetic energy) within the collision products is the azimuthal component (see Fig. 8), with the saturated magnetic field strength $B$ near the center $\simeq 10^{7}-10^{8}$ G, as shown for a representative case in Fig. 3. In general, the gradient of $B$, relative to the enclosed mass, is shallower for higher impact parameters ($b$) -- the surface $B$ is a few G for $b=0.1$, while it is $\simeq 10^{3}$ G for $b=0.5$. The mass-weighted average of magnetic flux is $\simeq 10^{26}-10^{28}$ G cm2 and surface magnetic flux of $10^{23}-10^{26}$ G cm2. If the magnetic flux is conserved, the surface magnetic field of the thermally relaxed collision product would be $10-10^{4}$. The dipole field ($\sqrt{(B^{r})^{2}+(B^{z})^{2}}$, where $B^{r}$ is the radial component and $B^{z}$ is the vertical component), more relevant for the magnetic braking and, thus, the spin down of the collision product, is consistently smaller than the total field by a factor of a few.
Figure 4: Density distribution of the products in collisions between a $0.7\,\mathrm{M}_{\odot}$ star and $0.68\,\mathrm{M}_{\odot}$ star with $b=0.25$ (top) and $0.5$ (bottom) along an $x-z$ slice passing through the center of mass, overlaid with equidensity lines (white contours). The density is measured at $t\simeq 45$ hours since the first contact. As $b$ increases, the overall shape of the product becomes more oblate because of larger angular momentum. Markedly, for $b=0.5$, a noticeably flattened structure along the mid-plane appears between the central part of the product and the outermost envelope, potentially indicative of disk formation.
@@
Figure 5: Mass-weighted average of the temperature $T_{\rm mid}$ in the mid-plane (top) and the ratio of the rotational velocity in the mid-plane to the Keplerian velocity $v_{\rm mid}^{\phi}/v^{\phi}_{\rm Kep}$ (bottom), as a function of distance from the center of mass of the collision products in collisions involving $M_{1}=0.7\,\mathrm{M}_{\odot}$, with different impact parameters $b$. The temperature profile shows a noticeable slope change at $0.7-0.8\,\mathrm{R}_{\odot}$ for $b\leq 0.25$ (except for one case with $M_{2}=0.4\,\mathrm{M}_{\odot}$ and $b=0.25$) and $1-2\,\mathrm{R}_{\odot}$ for $b=0.5$ (solid lines). The velocity ratio profiles show the similar features.